A right circular cone is inscribed in a cube having a diagonal which
measures 20 square root of 3 cm.

1. What is the altitude of the cone?
1
Expert's answer
2020-12-01T06:18:08-0500

The diagonal "d" of the cube with the side "a" is


"d^2=a^2+a^2+a^2=3a^2"

"d=\sqrt{3}a"

Given "d=20\sqrt{3}\ cm."

Then


"a=\dfrac{d}{\sqrt{3}}=\dfrac{20\sqrt{3}}{\sqrt{3}}=20(cm)"

The radius "r" of the circle inscribed in the square with the side "a" is


"r=\dfrac{a}{2}"

Then we have the right circular cone inscribed in a cube with the side "a"


"radius=r=\dfrac{a}{2}, \ height=h=a"

"h=20\ cm"

The altitude of the cone is 20 cm.